OpenStudy (anonymous):

A system of equations is shown below: 2x = 5y + 4 3x - 2y = -16 What is the solution to this system of equations? (-8, -4) (8, 4) (-4, -8) (4, 8)

OpenStudy (anonymous):

will medal :)

OpenStudy (texaschic101):

2x = 5y + 4 2x - 5y = 4 3x - 2y = -16 -->(-2) 2x - 5y = 4 ---->(3) ----------------- -6x + 4y = 32 (result of multiplying by -2) 6x - 15y = 12 (result of multiplying by 3) ----------------add -11y = 44 y = -4 2x = 5y + 4 2x = 5(-4) + 4 2x = -20 + 4 2x = -16 x = -8 solution is (-8,-4)

OpenStudy (anonymous):

AWESOME :)))) HELP ME WITH ONE MORE PLEASE? WILL MEDAL AGAIN <3 @texaschic101